OpenStudy (anonymous):

At what temperature would the most probale speed of CO2 molecules be twice that at 323K?

OpenStudy (anonymous):

MPV = sqrt 2RT/M

OpenStudy (anonymous):

@UnkleRhaukus plz help to figure out this

OpenStudy (anonymous):

@eigenschmeigen

OpenStudy (anonymous):

@ujjwal @UnkleRhaukus any idea

OpenStudy (unklerhaukus):

\[v_{\text{rms}}=\sqrt{\frac{3RT}{M}}\]

OpenStudy (anonymous):

most probale speed

OpenStudy (anonymous):

plzz read the quest once more

OpenStudy (unklerhaukus):

\[M_{\text{CO}_2}=44[\text{amu}]\]

OpenStudy (unklerhaukus):

http://www.physicsforums.com/showthread.php?t=274201

OpenStudy (ujjwal):

you just need to use the relation \[v=k\sqrt T\] at two different absolute temperatures. Here k is constant. \[v_1 =k\sqrt {323}\]\[v_2=k\sqrt {T_2}\]But \(v_2=2v_1\) so, we have \[\frac{v_1}{2v_1}=\sqrt{\frac{323}{T_2}}\]so, \(T_2\)= 1292 K=1019 degrees.

OpenStudy (anonymous):

yup u r correct

OpenStudy (anonymous):

Let me try to undersstand

OpenStudy (anonymous):

i got it thxxxx @ujjwal

OpenStudy (ujjwal):

wc. :)

OpenStudy (anonymous):

Plzz look out my next question! @ujjwal